OpenStudy (anonymous):

On Monday morning, Colin filled his car with six and one-half gallons of gas, then used three and one-fourth gallons driving. On Tuesday, he used up 2 gallons driving. On Tuesday night, his gas tank had four and one-fourth gallons remaining. How much gas was in Colin's tank before he filled up on Monday morning? 3 gal three and one-fourth gal three and one-half gal three and three-fourths gal

OpenStudy (texaschic101):

x + 6 1/2 - 3 1/4 - 2 = 4 1/4 (change everything to improper fractions) x + 13/2 - 13/4 - 2/1 = 17/4 (multiply everything by LCD, which is 4) 4x + 26 - 13 - 8 = 17 -- combine like terms 4x + 5 = 17 --- subtract 5 4x = 17 - 5 4x = 12 -- divide by 4 x = 3 gallons